OpenStudy (anonymous):

Solve the triangle. B = 36°, a = 41, c = 17

jagr2713 (jagr2713):

use a2+b2=c2

jagr2713 (jagr2713):

wait i think i am wrong

OpenStudy (jdoe0001):

have you covered the "Law of Cosines" yet?

OpenStudy (matt101):

You can't use the Pythagorean Theorem here because we don't know that this is a right-angled triangle. We have 2 sides and an angle, so I would use the Cosine Law.

jagr2713 (jagr2713):

yea i was thinking that to @matt101

OpenStudy (jdoe0001):

have you fainted yet? =) did you recover yet?

OpenStudy (jdoe0001):

well... I guess if you haven't... wouldn't be able to answer anyhow

OpenStudy (anonymous):

OS is lagging sorry. But yes I have covered the Laws of Cosine.

OpenStudy (jdoe0001):

\(\bf \textit{Law of Cosines}\\ \quad \\ b^2 = {\color{blue}{ a}}^2+{\color{red}{ c}}^2-(2{\color{blue}{ a}}{\color{red}{ c}})cos(B)\implies b = \sqrt{{\color{blue}{ a}}^2+{\color{red}{ c}}^2-(2{\color{blue}{ a}}{\color{red}{ c}})cos(B)}\) so that'd get you side "b" then you'd just need to find the other 2 angles

OpenStudy (anonymous):

Would it look something like this \[b^2=41^2+17^2-2(41)(17)CosB\]

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

so then once you have side "b", then you'd need to find the other angles since the sum of all internal angles in a triangle is 180 if you just find one of the other angles, you can pretty much get the other from 180-(sum of the other two angles)

OpenStudy (anonymous):

So then it would be \[b^2=1681+289-1394CosB\]\[b^2=1970-1394CosB\]\[b^2=576CosB\]Do I find the square root of 576?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

So b is 24? That isn't one of my answer choices. Did I do something wrong?

OpenStudy (jdoe0001):

so... once you found "b" you have all 3 sides, "a", "b" and "c" so, to find another angle just solve for cosine, that is \(\bf c^2=a^2+b^2-2ab\cdot cos(C)\implies c^2-a^2-b^2=-2ab\cdot cos(C) \\ \quad \\ \cfrac{c^2-a^2-b^2}{-2ab}= cos(C)\implies cos^{-1}\left[\cfrac{c^2-a^2-b^2}{-2ab}\right]=cos^{-1}[cos(C)] \\ \quad \\ cos^{-1}\left[\cfrac{{\color{red}{ c}}^2-{\color{blue}{ a}}^2-b^2}{-2{\color{blue}{ a}}b}\right]=\measuredangle C\)

OpenStudy (jdoe0001):

and angle A as you'd know, it'll be 180 - ( B+C)

OpenStudy (anonymous):

So C would be 20?

OpenStudy (jdoe0001):

20.14 approximately, in degrees, yes

OpenStudy (anonymous):

Awesome! So then A would be 124, if my calculations are correct.

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thank you for your help. You made this so easy to understand! :)

OpenStudy (jdoe0001):

yw